40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 13 19:36:36 2014, Koji, Summary, LSC, PRM 2nd/4th violin filter added 
    Reply  Tue Feb 25 18:16:22 2014, Jenne, Summary, LSC, Even more violin filters PRMvio2.png
    Reply  Tue Apr 1 21:45:01 2014, Jenne, Summary, LSC, PRM violin notch not at correct freq? 
Message ID: 9772     Entry time: Tue Apr 1 21:45:01 2014     In reply to: 9634
Author: Jenne 
Type: Summary 
Category: LSC 
Subject: PRM violin notch not at correct freq? 

Koji and I have the PRMI locked right now, and we hear a very strong violin mode ringing up, at 628Hz.  This is, according to Koji's elog 9634, what we expect the PRM's violin mode to be.  However, the current PRM violin mode notch is really more of a bandstop filter, between 622-670Hz.  At 628Hz, it has a suppression of about -20dB.  If I try to increase the width of this notch by making it 612-670Hz, the PRMI won't hold lock.

We're leaving this as a daytime task for tomorrow, since we're in the middle of taking data to show that Koji's new ASC filter design (slightly tuned from his elog 9769) works well.

Edit:  I have moved the PRM violin notch frequency over to 612-660 Hz, and after letting it sit for a while (while locked on PRMI), the violin mode has settled down.  Interestingly, if I compare the spectrum with and without the 1st order violin mode notch, it looks like the 2nd order mode changes from 1256Hz to 1303Hz.  I don't know what is going on here, but we already have notches for both of those frequencies.

ELOG V3.1.3-